# SonicSieve: Bringing Directional Speech Extraction to Smartphones Using Acoustic Microstructures

### v1

@@ ABSTRACT.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 38th Annual ACM Symposium on User Interface Software and Technology; Sep. 28–Oct. 01, 2025; Busan, Korea
@@ 1. INTRODUCTION
Here, we ask the question: Can we enable directional speech extraction on smartphones? A positive answer can enable smartphones to intelligently amplify target speakers and suppress unwanted interference. Imagine having a meeting in a noisy restaurant, and you only want to capture the speech of colleagues at your table for clear meeting notes and to remotely stream to other colleagues back at the office (Fig. 1). Or imagine being a student in a classroom and you want to record the sound of the lecturer without needing a wearable microphone or a dedicated setup.
@@
We present SonicSieve, the first intelligent directional speech extraction system on smartphones that uses a bio-inspired acoustic microstructure to embed distinct directional cues into incoming speech sounds from different angles. Our design is inspired by the directional hearing mechanisms found in nature. While human ears function like a two-microphone array, they are able to achieve directional hearing by leveraging spatial cues from the interaction of sound with the structure of the head and pinna (Brughera et al., 2019). We build on this idea and design a lightweight, miniature 3D-printed cylinder patterned with coded holes that are designed to diffract incoming speech, and create a set of virtual sound sources. The resulting interference patterns form a distinct acoustic signature that changes depending on the direction of a sound, effectively embedding spatial information into the signal.
Our microstructure attaches to the in-line microphone of low-cost wired earbuds, which can be plugged into a smartphone. The system captures audio mixtures using both the in-line microphone and the smartphone’s co-located built-in top microphone, which serves as a reference (Fig. 1) and enables the system to function reliably across diverse sound sources. A real-time neural network running on the smartphone processes these audio mixtures, allowing users to directionally filter speech from different angular sectors. We also design a smartphone user interface that lets users divide the surrounding space into a semicircular region with six 30° sectors. Users can select one or multiple sectors simultaneously, enabling functionality in multi-speaker scenarios (Fig. 13) such as transcribing presentations with multiple speakers or conducting remote meetings in noisy environments like a park.
@@ 2. RELATED WORK
Beamforming using statistical algorithms. Traditional beamforming techniques are typically based on statistical algorithms, including non-adaptive approaches such as Bartlett (delay-and-sum) and superdirective beamformers, as well as adaptive methods like minimum-variant distortionless-response (MVDR), and linearly constrained minimum variance (LCMV). However, these algorithms are designed for conventional microphone arrays with fixed and often uniform geometries, and do not account for the complex direction-dependent transformations introduced by acoustic microstructures. As a result, they cannot be directly applied to microstructure-enabled systems like ours, where spatial cues arise not only from microphone placement but also from the physical design of the structure. Moreover, statistical beamformers are limited to capturing spatial cues, and lack the ability to model acoustic cues as neural beamformers do.
@@
Spatial sensing on smartphones. Prior work has focused on estimating DoA (Küçük et al., 2019; Zhou et al., 2024; Tokgöz et al., 2020), and beamforming (Gaubitch et al., 2014) on smartphones using statistical algorithms and neural networks. However, the performance of these systems is fundamentally limited by the limited number of microphones. Ad hoc microphone arrays composed of smartphones have also been proposed for spatial sensing. These systems emulate larger arrays by synchronizing multiple devices. However, they rely on complex synchronization protocols (Sur et al., 2014) such as aligning smartphone CPU clocks using WiFi beacon timestamps, synchronizing audio I/O clocks through a backend server, or using time-of-arrival techniques (Li et al., 2017; Peng et al., 2007), which can make deployment in real-world scenarios challenging.
@@ 3. SONIC SIEVE
Our system has three main components which we describe below:
  • Acoustic microstructure for speech on smartphones. We present a compact acoustic microstructure that enables directional speech filtering by creating spatial cues detectable by microphones. Our design attaches to the in-line microphone of low-cost wired earbuds which can be plugged into smartphones. It is easily fabricated using standard 3D printing techniques and generalizes across different sound signals and environments. Our design is optimized for spatial diversity while remaining pocket-sized.
  • Real-time directional speech extraction network. A directional hearing algorithm is required to extract and interpret the spatial cues from the microstructure. A challenge however with conventional beamformers using signal processing or neural networks is that they are typically designed for uniform microphone array geometries. In contrast, our microstructure introduces spatial diversity through holes that act as virtual sound sources, and produces a complex multipath profile that deviates from these assumptions. We design a real-time neural network that learns the spatial and acoustic cues introduced by the microstructure to simultaneously extract speech from multiple selected directions. This enables multi-speaker transcription in scenarios such as interviews and group discussions.
3.1. Speech-aware acoustic microstructure for smartphones
@@
(a) Without Microstructure
(b) With Microstructure
Figure 3. Frequency response ($M_{\theta}(f)$) at the microphone without and with microstructure for incoming sound signals at different angles of arrival. The plot shows that the microstructure creates greater variation across different angles, which can be leveraged for more effective directional hearing.
[…]
In single-microphone devices, determining the direction of a sound source is difficult as recordings from different angles are similar (Saxena and Ng, 2009). While reflections caused by reflectors in the room can encode information about the sound source direction, these cues are specific to one environment and do not generalize (An et al., 2018, 2022). Microphone arrays can estimate the direction of a sound source by analyzing differences in time and amplitude across multiple microphones, with accuracy improving as microphone count increases (Chung et al., 2022; Berg et al., 2024).
@@
Interestingly, while traditional two-microphone arrays can only localize along one axis with spatial ambiguities (Desai and Mehendale, 2022; Francl and McDermott, 2022), humans can localize sound effectively with just two ears (Van Opstal, 2016). This is due to the geometry of the head and ears which create distinctive reflecting and scattering effects that vary with sound angle (Xie, 2013). Such a principle is common in nature’s directional hearing system, like asymmetric owl ears (Coles and Guppy, 1988). This capability stems from the complex geometry of the head and outer ears, which creates distinctive reflecting and scattering effects that vary with sound angle (Xie, 2013). This principle appears throughout nature’s directional hearing systems, like the asymmetric ear structures of owls that enable their exceptional sound localization abilities (Coles and Guppy, 1988). We take inspiration from this phenomena and aim to create an acoustic microstructure that embeds direction-dependent spatial cues through similar physical effects.
3.1.1. Microstructure Design.
We build on top of prior designs (Garg et al., 2021) consisting of a hollowed-out cylinder that encases a microphone. Below we describe our methodology for selecting key design properties of our microstructure.
  1. (1)
    Surface holes with $\sim 1~{}mm$ diameter on the exterior are smaller than speech sound wavelengths, causing incoming sounds to bend (diffract) and create complex direction-dependent sound patterns inside the structure (Pierce, 1974).
  2. (2)
    Capillary tubes connect these holes to the interior, with varying lengths and shapes that modify how sound waves travel through them (Peat, 1994), creating unique phase patterns for sounds from different directions (Tijdeman, 1975).
  3. (3)
    Helmholtz resonators are small spheres placed along the tubes that act like tiny musical instruments. Each resonator selectively amplifies different frequency bands (Tang and Sirignano, 1973; Yuan et al., 2024b), and enhances the spatial diversity of the resulting acoustic signatures.
@@
Specifically, for a sound source arriving at the microstructure from a relative angle $\theta$ and propagating through the internal structure, the microstructure applies a direction-specific filter $M_{\theta}(f)$ to the sound where $f$ is the frequency (Fig. 2). We show in Fig. 3(b) the direction-specific filter $M_{\theta}(f)$ at four different angles without and with the microstructure. The plot shows that the microstructure introduces substantially greater variation across different directions of arrival, and these spatial cues can be used to build a better directional speech extraction network.
Given the microstructure design, our next goal is to estimate such a directional signature $M_{\theta}(f)$ from the received signal and leverage it to enhance the speech extraction network.
Generalizing across sound sources and environments. In our discussion thus far, we have focused on how the microstructure embeds direction-specific cues onto the signal. In practice, the signal at the microphone is also affected by two other factors: the sound source $S(f)$ and the multipath environment effect $H(f)$ (Garg et al., 2021; Yuan et al., 2025), and the received signal at the microphone combines these effects as:
(1) $R(f)=S(f)H(f)M_{\theta}(f)$
If the source $S(f)$ is pre-defined and the environmental reflection $H(f)$ is measured beforehand, we can obtain the microstructure’s direction filter by simply canceling their effects through division in the frequency domain (i.e. $M_{\theta}(f)=\frac{R(f)}{S(f)H(f)}$). However, in practice we cannot know the source signal or the environmental reflection.
Figure 4. Effect of separation between earphone’s in-line mic and phone’s top mic on spectral variation. We compute the ratio of the received signals at the two microphones ($H_{ratio}(f)$) and estimate the variation across environments (lower is better). When the mics are closely spaced, there is less spectral variation.
[…]
@@
To eliminate such source and environment dependency, we incorporate a secondary reference microphone placed outside the microstructure (the built-in microphone on the smartphone). The reference microphone records the sound unmodified by the microstructure simultaneously. The received signal can be denoted as:
(2) $R_{ref}(f)=S(f)H^{\prime}(f)$
We note that the corresponding environment reflection $H^{\prime}(f)$ differs slightly from $H(f)$ as the reference microphone is not positioned at exactly the same location as the microphone inside the microstructure. To isolate the effect of microstructure, we divide the signal received by the main inside microphone and the reference microphone:
(3) $\frac{R(f)}{R_{ref}(f)}=\frac{H(f)}{H^{\prime}(f)}M_{\theta}(f)=H_{ratio}(f)M_{\theta}(f)$
In this way, although the source effect is eliminated, the impact of environmental factors $H_{ratio}$ is not completely removed. To minimize the residual effect of the environmental reflections, we position the reference microphone as close as possible to the microstructure. This proximity minimizes path differences between the two microphones, ensuring that sound waves encounter similar environmental reflections before reaching either microphone, thereby reducing the variation in $H_{ratio}(f)$.
We investigate this empirically and perform a benchmark experiment by placing the in-line mic of the earphone and the built-in top mic of the smartphone at different separation distances and measure the $H_{ratio}(f)$ in 160 locations in total across 4 different environments. Our results shown in Fig. 4 demonstrate that the amplitude variation of $H_{ratio}(f)$ increases with microphone separation. At 1 cm separation, variations remain below 2.2 dB across frequencies, while at 16 cm, variations exceed 6 dB at higher frequencies. Based on this information, our prototype is designed so that the in-line microphone is within 1 cm of the smartphone’s built-in top mic.
3.1.2. Optimizing the microstructure design.
Next, we detail key design parameters of the microstructure and their impact on system performance. We then outline our approach for selecting these parameters in the final design. Our goal in optimizing the microstructure is to maximize spatial diversity, in other words, the variation in frequency responses for sound arriving from different directions. To do this, we design a structure that creates unique directional acoustic signatures through three distinct structural elements:
Material Selection. Our goal in selecting a material for the microstructure is to maximize attenuation of incoming sound, ensuring that sound primarily travels through the holes and diffracts as a virtual sound source rather than passing through the walls. To evaluate this, we transmit a 1–8 kHz chirp, covering the frequency range important for understanding human speech (Kunchur, 2023)—at a sound level of 70 dBA which is on the order of normal conversation. We measure and plot the frequency response at the microphone in three scenarios: without the microstructure, with a silicone microstructure, and with a resin microstructure (Fig. 5). The resin microstructure, which we use in our final design, achieves an average sound reduction of 33 dB across the tested frequencies, outperforming the silicone microstructure, which achieves a reduction of 17 dB.
Microstructure diameter. We measure how the diameter of the microstructure affects its spatial diversity. We measure the overall spatial diversity by calculating $\left\lVert M_{\theta_{1}}(f)-M_{\theta_{2}}(f)\right\rVert_{2}$) for angle range from 0° to 180° with 1° resolution for all pairwise angles. We varied the microstructure diameter across 10, 15, and 20 mm while measuring spatial diversity, as shown in Fig. 6 (b–d). The plot shows that increasing the diameter leads to greater spatial diversity. However, this improvement comes with a tradeoff: larger microstructures can obstruct the smartphone screen, potentially hindering usability. In our design, we selected a 20 mm diameter as a balance between increased spatial diversity and user experience.
Number of surface holes. We examine the effect of the number of surface holes on spatial diversity. With only a single hole, directional sensitivity is limited as sounds arriving from directions other than the hole are largely attenuated by the microstructure and cannot be reliably distinguished. On the other hand, if the structure has too many holes, it behaves similarly to having no microstructure at all, with little spatial diversity. We compare the spatial diversity achieved with 6 holes (Fig. 6e) versus 10 holes (Fig. 6d). Empirically, we observe that the configuration with 6 holes provides greater spatial diversity than the 10-hole design, and we select 6 holes for our final design.

### v2

@@ ABSTRACT.
conference: ACM Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spaincopyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXX
@@ 1. INTRODUCTION
In this paper, we ask the question: Can we enable directional speech extraction on smartphones? A positive answer can enable smartphones to intelligently amplify target speakers and suppress unwanted interference. Imagine having a meeting in a noisy restaurant, and you only want to capture the speech of colleagues at your table for clear meeting notes and to remotely stream to other colleagues back at the office (Fig. 1). Or imagine being a student in a classroom and you want to record the sound of the lecturer without needing a wearable microphone or a dedicated setup.
@@
We present SonicSieve, the first intelligent directional speech extraction system on smartphones that uses a bio-inspired acoustic microstructure to embed distinct directional cues into incoming speech sounds from different angles. The design of the acoustic microstructure is inspired by directional hearing mechanisms found in nature. While human ears function like a two-microphone array, they are able to achieve directional hearing by leveraging spatial cues from the interaction of sound with the structure of the head and pinna (Brughera et al., 2019). Perhaps closest to SonicSieve in prior art is Owlet (Garg et al., 2021), a lightweight, miniature 3D-printed cylinder patterned with coded holes to encode spatial information into the signals arriving from different directions, enabling Direction-of-Arrival (DoA) estimation for a single sound source. Inspired by Owlet, SonicSieve address a real-world user-centered challenge: going beyond localizing where a sound signal originates, SonicSieve tackles directional speech extraction that determines what to extract from an acoustic mixture in a noisy multi-speaker environment.
To this end, we co-design both novel hardware and software solutions for SonicSieve. First, we optimize the dimensions and the coded pattern of the microstructure specifically for speech frequencies. We then create a practical design that attaches the microstructure to the in-line microphone of low-cost wired earbuds, which can be plugged into a smartphone. The system captures audio mixtures using both the in-line microphone and the smartphone’s co-located built-in top microphone, which serves as a reference (Fig. 1) and enables the system to function reliably across diverse sound sources. Second, we develop an end-to-end neural network that runs in real-time on the smartphone, processing the audio from both microphones to interpret the rich spatial cues and separate the speech from the target directions. We also design a smartphone user interface that lets users divide the surrounding space into a semicircular region with six 30° sectors. Users can select one or multiple sectors simultaneously, enabling functionality in multi-speaker scenarios (Fig. 15) such as transcribing presentations with multiple speakers or conducting remote meetings in noisy environments like a park.
@@ 2. RELATED WORK
Reference Mics Single device Directional speech extraction Smartphone compatible Real-world evaluation Real-time
Conventional mic array 4 to 8+ N/A N/A
Dia (2014) (Sur et al., 2014) 2 to 8
Owlet (2021) (Garg et al., 2021) 2
DSENet (2023) (Kovalyov et al., 2023) 3
EarCase (2023) (Li et al., 2023) 2 DNS
Pandey et al. (2024) (Pandey et al., 2024) 8
SonicSieve (ours) 2
Table 1. Comparison of SonicSieve with related systems. SonicSieve enables directional speech extraction across six $30^{\circ}$ sectors on smartphones in real-time, with evaluation in real-world environments. (DNS = did not specify)
@@
Beamforming using statistical algorithms. Traditional beamforming techniques are typically based on statistical algorithms, including non-adaptive approaches such as Bartlett (delay-and-sum) and superdirective beamformers, as well as adaptive methods like minimum-variant distortionless-response (MVDR), and linearly constrained minimum variance (LCMV). However, these algorithms are designed for conventional microphone arrays with fixed geometries, and do not account for the complex direction-dependent transformations introduced by acoustic microstructures. As a result, they cannot be directly applied to microstructure-enabled systems like ours, where spatial cues arise not only from microphone placement but also from the physical design of the structure. Moreover, statistical beamformers are limited to capturing spatial cues, and lack the ability to model acoustic cues as neural beamformers do.
@@
Spatial sensing on smartphones. Prior work has focused on estimating DoA (Küçük et al., 2019; Zhou et al., 2024; Tokgöz et al., 2020), and beamforming (Gaubitch et al., 2014) on smartphones using statistical algorithms and neural networks. However, the performance of these systems is fundamentally limited by the number of microphones. Ad hoc microphone arrays composed of smartphones have also been proposed for spatial sensing (Sur et al., 2014). These systems emulate larger arrays by synchronizing multiple devices. However, they rely on complex synchronization protocols such as aligning smartphone CPU clocks using WiFi beacon timestamps, synchronizing audio I/O clocks through a backend server, or using time-of-arrival techniques (Li et al., 2017a; Peng et al., 2007), which can make deployment in real-world scenarios challenging.
@@ 3. SONIC SIEVE
This section outlines the technical approach of SonicSieve. We start with a high-level overview of how our system uses an acoustic microstructure for directional speech extraction. Then, we will detail the techniques for optimizing this microstructure, integrating it with a smartphone, using a neural network to extract speech in real-time, and real-world dataset collections.
(a) Without Microstructure
(b) With Microstructure
Figure 2. Effect of microstructure on incoming sound signals across different angles of arrival. The microstructure introduces larger variations in the frequency response $M_{\theta}(f)$ across different angles, providing enhanced spatial cues that can be leveraged for directional speech extraction.
[…]
@@
3.1. System Overview
Traditional directional speech extraction relies on microphone arrays that use beamforming to isolate sound from specific directions. These systems work by analyzing timing and amplitude differences across multiple microphones—when sound arrives from a particular direction $\theta$, each microphone in the array receives the signal at slightly different times and intensities due to their spatial arrangement. The system can be modeled as:
(1) $\mathbf{x}(f,t)=\mathbf{a}(f,\theta)s_{\theta}(f,t)+\mathbf{u}(f,t)$
where $\mathbf{a}(f,\theta)$ encodes the directional information based on known geometric delays between microphone channels, and $\mathbf{u}(f,t)$ represents interference and noise. Beamforming algorithms then apply spatial filtering to enhance signals from the target direction while suppressing others.
While various beamforming techniques work reliably for directional audio enhancement (Adel et al., 2012; Araki et al., 2007), they typically requires a microphone array with multiple channels (usually 4-8) to handle the spatial ambiguity (Desai and Mehendale, 2022; Francl and McDermott, 2022) and achieve satisfactory performance, which does not exist on smartphones.
@@ 3.1. System Overview
Instead of encoding the spatial information through microphone array geometry, SonicSieve leverages a passive acoustic microstructure that directly embeds direction-dependent spatial cues into the signal. Such a design is inspired by nature’s directional hearing system. For example, humans can localize and separated sound effectively with just two ears, which is due to the geometry of the head and ears creating distinctive reflecting and scattering effects that vary with sound angle (Van Opstal, 2016; Xie, 2013). Similarly, owls achieve exceptional sound localization through asymmetric ear structures (Coles and Guppy, 1988).
@@
Specifically, our acoustic microstructure applies a direction-specific filter $M_{\theta}(f)$ to a sound source arriving at the microstructure from a relative angle $\theta$ and propagating through the internal structure. We present an example in Fig. 2(b) of the direction-specific filter $M_{\theta}(f)$ at four different angles without and with the microstructure. The plot shows that the microstructure can introduce substantially greater variation across different directions of arrival.
The received signal at the microphone inside the microstructure can be denoted as:
(2) $x(f,t)=M_{\theta}(f)s(f,t)+u(f,t)$
However, with only single-channel audio, we can not isolate the spatial filter $M_{\theta}(f)$ from the signal source $s(f,t)$. Thus, we incorporate a secondary reference microphone placed outside the microstructure. The reference microphone records the sound unmodified by the microstructure simultaneously:
(3) $x_{ref}(f,t)=s^{\prime}(f,t)+u_{ref}(f,t)$
We note that the received signal at the reference microphone from the target source $s^{\prime}(f,t)$ differs slightly from the one at the inside microphone $s(f,t)$, since the two microphones are not positioned at the exact same location, such that the arriving signals are propagated through slightly different environmental reflections. To enable robust estimation of the spatial cues across different environments, the two microphones need to be positioned as close as possible.
Finally, we develop a deep neural network that processes the two-channel input to extract directional speech from the desired direction. Specifically, our directional speech extraction network $h$ estimates the target speech from the selected directions $\mathbf{\theta}$:
(4) $\hat{s}_{\mathbf{\theta}}(f,t)=h\left(x(f,t),x_{ref}(f,t),\mathbf{\theta}\right)$
To summarize, our system has four main components, which we describe below and present the details in the following sections:
  • Acoustic microstructure optimized for speech. We present a compact acoustic microstructure that encodes spatial cues into the signal detectable by microphones. Building on prior microstructure work (Garg et al., 2021), we further optimize the design specifically for spatial diversity at speech frequencies.
  • Integration of acoustic microstructure to smartphones. We present our design that can attach the acoustic structure to the in-line microphone of low-cost wired earbuds, which can be plugged into smartphones. This approach offers two key benefits: the in-line mic can be positioned close to the built-in mic to minimize signal differences, and the design can potentially be generalized to different smartphone models.
  • Real-time directional speech extraction network. While conventional beamformers work with known array geometries, they cannot handle the complex propagation patterns created by microstructures. We design a real-time neural network that learns the spatial and acoustic cues from our microstructure to extract speech from multiple selected directions simultaneously, enabling multi-speaker transcription for interviews and group discussions.
3.2. Speech-aware acoustic microstructure
We build upon Owlet’s microstructure design (Garg et al., 2021) which consists of a hollowed-out cylinder that encases a microphone. The design incorporates three key structural elements: surface holes that act as virtual sound sources, capillary tubes with varying lengths that modify sound wave propagation, and Helmholtz resonators that selectively amplify different frequency bands. Together, these elements create unique acoustic signatures for sounds arriving from different angles. As visualized in Fig. 3, when sound arrives from angle $\theta$, the microstructure applies a direction-specific filter $M_{\theta}(f)$. This directional fingerprint enables our system to distinguish between sounds from different locations.
In this section, we present our approach to optimize the acoustic microstructure and detail the key design factors that impact the overall system performance below:
@@ 3.2. Speech-aware acoustic microstructure
Figure 5. Effect of microstructure diameter on spatial diversity. The microstructure with a larger diameter (20 mm) provides an overall higher spatial diversity across the speech frequencies.
[…]
The figure illustrates the effect of different microstructure diameters on spatial diversity across a range of frequencies. The x-axis is labeled ”Frequency(Hz)” and ranges from 0 to 8000 Hz. The y-axis is labeled ”Spatial Diversity” and ranges from 0 to 8. The figure contains four lines, each representing a different condition. The black solid line represents a condition ”w/o Microstructure”(without microstructure). The red dashed line represents a 15 mm microstructure diameter, labeled as ”15 mm(Owlet)”. The blue dashed represents a 18 mm microstructure diameter. The green dotted line represents a 20 mm microstructure diameter. The plot shows that all four lines generally increase in spatial diversity as frequency increases, but their behavior varies. The ”w/o Microstructure” line remains at the lowest spatial diversity across all frequencies. The 15 mm line shows a slight increase but remains relatively low. The 18 mm and 20 mm lines show significantly higher spatial diversity, with the 20 mm line generally outperforming the 18 mm line. The microstructure with a larger diameter provides an overall higher spatial diversity across the speech frequencies.
Material Selection for Acoustic Attenuation. Our material selection aims to maximize sound attenuation, ensuring audio travels through the designed holes rather than leaking through the walls. We tested three configurations using a 1–8 kHz chirp at 70 dBA (normal conversation level): no microstructure, silicone microstructure, and resin microstructure.
As shown in Fig. 5, the resin microstructure achieves 33 dB average sound reduction across tested frequencies, significantly outperforming silicone (17 dB reduction). This superior attenuation ensures sound primarily diffracts through the holes as intended, making resin our material of choice.
Spatial Diversity Optimization. To optimize performance for directional speech extraction, we maximize spatial diversity—the variation in how the structure responds to speech from different directions. Mathematically, we seek to maximize the variance $\mathbb{V}_{\theta}\left[|M_{\theta}(f)|\right]$, ensuring speech signals from different angles produce maximally different frequency responses. Speech signals are wideband acoustic signals with frequency components mainly under 8 kHz, and the components that are most critical for intelligibility are in 1-4 kHz (Kunchur, 2023). We conduct experiments using wideband audio under 8 kHz in a constrained environment (anechoic chamber) on different microstructure designs.
To measure spatial diversity, we record speech signals from 0° to 180° at 1° resolution using speakers and microphones inside and outside the microstructure. Since the measurements are taken in an anechoic chamber, we can assume the environmental reflections and background noise are negligible (i.e. $s^{\prime}(f,t)=s(f,t)$ and $u(f,t)=0$). From Eq. 2 and Eq. 3, the spatial diversity can then be calculated by:
(5) $\mathbb{V}_{\theta}\left[|M_{\theta}(f)|\right]=\mathbb{V}_{\theta}\left[\frac{|x(f)|}{|x_{ref}(f)|}\right]$
where $x(f)$ and $x_{ref}(f)$ are signals from the internal and reference microphones. Based on the measurement setup, we optimize the following two key factors to improve the spatial diversity.
  • Microstructure Diameter. Fig. 5 shows that Owlet’s original 15 mm design achieves spatial diversity mainly at 7–8 kHz, with poor performance below 7 kHz, which can be potentially problematic for speech signals. Based on the principle that microstructure size correlates with resonating wavelengths, we conduct measurements on larger diameters. Both 18 mm and 20 mm diameters show improved diversity below 8 kHz, with 20 mm providing a better performance across speech frequencies. We further visualize spatial diversity across different angles by calculating the pairwise spectral distance (i.e. $\left\lVert M_{\theta_{1}}(f)-M_{\theta_{2}}(f)\right\rVert_{2}$) for the microstructure with diameters of 10 mm, 15 mm, and 20 mm. As shown in Fig. 6 (a-d), the microstructure with a higher diameter achieves higher spatial diversity across a wider range of angles. We use 20 mm as our final design. We note that we did not further increase the diameter as a larger microstructure may obstruct the smartphone screen and hinder usability.
@@
  • Number of Surface Holes. As we observed in Fig. 6(d), the spatial diversity across the angles is not evenly distributed—angles >60>60° show lower diversity than 0–60°. To address this, we further optimized the surface hole pattern. We empirically find that, too few holes limit directional sensitivity (sounds from non-hole directions get attenuated), while too many holes eliminate spatial diversity entirely. Thus, we selectively reduce the original ten-hole design to six holes by removing four holes in the 60–180° range. Fig. 6(e) demonstrates that this six-hole configuration outperforms the ten-hole design in overall spatial diversity. Therefore, our final design uses a 20 mm diameter cylinder with six strategically placed surface holes.
3.3. Integration of acoustic microstructure to smartphones
Figure 7. Illustration of the process to integrate acoustic microstructure to the smartphone. (a) UI on the smartphone app instructs the user to place the microstructure in the proper location. (b) Place the inline microphone of a wired earphone on the bottom holder of the microstructure. (c) The top part housing the optimized microstructure is attached to the holder using tiny screws. Proper alignment is achieved with the built-in top microphone of the smartphone.
[…]
The three sub-figures illustrate the process of integrate acoustic microstructure to the smartphone. The sub-figure (a) is a photo of a smartphone app user interface. The screen displays the name ”SonicSieve” at the bottom. A fan-like multicolored graphical element is positioned in the center, and a large circular button is at the bottom. At the top of the screen, there is a text prompt that says, ”Place Microstructure Here,” with a dashed rectangular box indicating where the microstructure should be placed. The sub-figures (b) and (c) are close-up photographs showing the physical assembly of the microstructure with a smartphone. The sub-figure (b) shows a black plastic holder attached to the bottom of the smartphone. An arrow points to a component labeled ”Inline Mic,” which is part of a wired earphone. The sub-figure (c) shows the full assembly from a different angle, revealing a top gray plastic part labeled ”Top Mic,” which houses the optimized microstructure. This top part is attached to the bottom holder using small screws. The figure as a whole visually explains the steps a user would take to correctly attach the acoustic microstructure to a smartphone for the system to function.
Figure 8. Effect of separation between earphone’s in-line mic and phone’s top mic on spectral variation. We compute the ratio of the received signals at the two microphones ($s_{ratio}(f)$) and estimate the variation across environments (lower is better). When the mics are closely spaced, there is less spectral variation.
[…]
Microstructure Fabrication. We implement our optimized microstructure design using computer-aided design and 3D printing. The microstructure geometry is modeled in Rhino 8111https://www.rhino3d.com/ and fabricated using stereolithography 3D printing. We use ELEGOO Standard Photopolymer Grey Resin [^2] https://us.elegoo.com/products/elegoo-standard-resin printed on a Formlabs Form 3 printer [^3] https://formlabs.com/3d-printers/form-3/, followed by post-processing with Form Wash for cleaning and Form Cure for curing. This fabrication process ensures precise hole dimensions and smooth internal surfaces critical for acoustic performance.
Assembly Design. Our integration approach attaches the microstructure to the in-line microphone of standard wired earphones, enabling compatibility with any smartphone equipped with an audio jack. As illustrated in Fig. 7, the system consists of two main components: a bottom holder that secures the earphone’s in-line microphone, and a top housing containing the optimized microstructure. The smartphone app guides users through proper placement, ensuring correct alignment between the microstructure and the device’s built-in top microphone.
The final assembly process involves three steps: (1) positioning the earphone’s in-line microphone in the bottom holder, (2) attaching the microstructure housing to the holder using precision-fitted connections, and (3) aligning the complete assembly with the smartphone’s top microphone. A silicone washer at the interface between components minimizes acoustic leakage and ensures consistent positioning.
Design Benefits. This integration approach offers two key advantages over device-specific solutions. First, it enables close microphone spacing—our design positions the in-line microphone within 1 cm of the smartphone’s built-in microphone, which is critical for minimizing environmental variation effects as established in Section 3.1. Second, the design generalizes across different smartphone models since it relies on standard audio jacks and commonly positioned top microphones rather than custom cases for specific devices.
Microphone Separation Validation. To validate the importance of close microphone spacing, we empirically measure how separation distance affects signal consistency across environments. Specifically, we measure the signal ratio of two microphones without attaching the microstructure, defined by:
(6) $s_{ratio}(f)=\frac{s(f)}{s_{ref}(f)}=\frac{x(f)}{x_{ref}(f)}$
where $x(f)$ and $x_{ref}(f)$ are signals from the inline and reference microphones.
We conducted experiments across 160 locations in 4 different environments, measuring the magnitude of $s_{ratio}(f)$ at separation distances of 1 cm, 8 cm, and 16 cm. Fig. 8 demonstrates that amplitude variation increases dramatically with microphone separation. At 1 cm separation, variations remain below 2.2 dB across all frequencies, while at 16 cm separation, variations exceed 6 dB at higher frequencies. This validates our design requirement for close microphone positioning and confirms that our 1 cm separation distance minimizes environmental artifacts that could interfere with directional cue extraction.
3.4. Directional speech extraction network
As described in Section 3.1, our microstructure encodes direction-specific spatial cues into the recorded audio signal. However, real-world acoustic environments contain multiple sound sources, environmental reflections, and noise that complicate the extraction of these spatial cues. We develop a deep neural network that processes the two-channel recordings (from the microstructure and reference microphones) to perform directional speech separation in real-time.
Our network design adapts a state-of-the-art real-time sound extraction architecture (Chen et al., 2024) for angular-based directional speech extraction. The system processes short audio chunks (8 ms) to enable real-time applications such as live audio streaming and meeting transcription on smartphones.
Pre-processing Module. Our system processes 2-channel audio input $\mathbf{x}=[x_{1},x_{2}]^{T}\in\mathbb{R}^{2\times T}$, where $x_{1}$ and $x_{2}$ correspond to the reference microphone and microstructure microphone signals respectively, with $T$ representing the audio sample length. To accommodate model lookahead requirements, we apply zero padding to create $\mathbf{x}\in\mathbb{R}^{2\times(T+\sigma)}$, where $\sigma$ is the lookahead sample count.
We implement a learnable Short-Time Fourier Transform (STFT) encoder with sufficiently large window size to enhance model stability across different reverberation environments (Cord-Landwehr et al., 2022). After applying the STFT encoder with discrete Fourier transform length $L$ and step size $\delta$, each microphone channel $i$ produces $X_{i}[f,t]\in\mathbb{R}^{F\times\frac{(T+\sigma-L)}{\delta}}$, where $F$ represents the number of frequency bins.